A sales aptitude test is administered to a national sample of sales professionals in the real estate industry. Nationally, the mean score on the test is 48, with a standard deviation of 12. Analysis of the sample's actual sales performance compared to the test results indicates that people with test scores above the 75th percentile tend to be outstanding salespeople. What test score represents the 75th percentile? (Round to the nearest whole number.)
Answer #1

Solution:

X = Mean + Z*SD

Mean = 48

SD = 12

Z for 75th percentile = 0.67448975 (by using z-table or excel)

X = 48 + 0.67448975*12

X = 56.093877

Answer: 56